# Gaze-LLE: Gaze Target Estimation via Large-Scale Learned Encoders

### v1

@@ 9 Runtime Analysis
Our ViT-Base model runs in 15ms ($\approx 66$fps) on a single NVIDIA RTX 4090 GPU. We compare the inference time of our model with existing methods in Fig 7(a). For Miao et al. [42], we include the auxiliary depth estimation model (DPT-Hybrid[49]) in runtime calculation. Compared to Miao et al., our approach is both faster, and achieves better performance. In fact, the inference time of the DPT-Hybrid depth model (17ms) exceeds the entire inference time of our approach. This result highlights the benefit of using a single encoder, both in inference speed and performance. Chong et al.’s approach [9], which does not use any models for auxiliary modalities like depth, runs faster than our model. However, this comes with a significant drop in performance compared to our method. As shown in Table 3, recent convolutional methods all use at least one auxiliary model to augment performance. While these approaches may use faster backbones than a ViT, requiring auxiliary models ultimately increases runtime.
@@
We also investigate how our model’s runtime scales with estimating the gaze for multiple people per image (Fig. 7(b)). We measure the inference time for 1-10 people per image for both our default method, and our variant that uses a head position token $(t_{\text{pos}})$ and decodes gaze via cross attention and a dot product with the scene features (Table 7 configuration 1b 2b). Because the majority of our model’s computation can be attributed to the DINOv2 scene encoder ($>$95% of computation), which is run once regardless of the number of people, our model’s runtime does not increase much with addition of more people (15ms for 1 person vs. 19ms for 10 people). The token variant of our model with cross attention scales even better, as it decodes gaze for all people from the same final feature map. However, as shown in Fig. 7(a), this is accompanied by a slight performance decrease.
@@ 10 Performance with Estimated Head Bounding Boxes
Tu et al. [64] report that 2-stream methods suffer major performance drops when using head bounding boxes from a detector rather than the dataset ground truth. In contrast, we observe almost no performance degradation when pairing our method with a YOLOv5 head detector trained on CrowdHuman [53, 1] (Tab. 15). This result demonstrates that our single-stream design, which uses a coarse, downsampled head position map, is less dependent on an exact head crop, and works well with out of the box head detections.
@@ 13 LoRA Backbones
To investigate if training the backbone improves performance, we explore using Low Rank Adaptation (LoRA) [26] on GazeFollow in Table 19. LoRA updates the backbone while introducing limited additional learnable parameters by learning weight update matrices as low rank decompositions. We update the query and value projections of the DINOv2 backbone using rank 16. We observe limited improvements, which we attribute to (1) the effectiveness of the frozen encoder’s feature representation for our task and (2) that our models with frozen encoders already achieve extremely close performance to the inter-rater performance of the human annotators, which serves as a soft upper bound on the L2 metrics.

### v2

@@ 9 Runtime Analysis
Our ViT-Base model runs in 15ms ($\approx 66$fps) on a single NVIDIA RTX 4090 GPU. We compare the inference time of our model with existing methods in Fig 7(a). For Miao et al. [43], we include the auxiliary depth estimation model (DPT-Hybrid[50]) in runtime calculation. Compared to Miao et al., our approach is both faster, and achieves better performance. In fact, the inference time of the DPT-Hybrid depth model (17ms) exceeds the entire inference time of our approach. This result highlights the benefit of using a single encoder, both in inference speed and performance. Chong et al.’s approach [9], which does not use any models for auxiliary modalities like depth, runs faster than our model. However, this comes with a significant drop in performance compared to our method. As shown in Tab. 3, recent convolutional methods all use at least one auxiliary model to augment performance. While these approaches may use faster backbones than a ViT, requiring auxiliary models ultimately increases runtime.
@@
We also investigate how our model’s runtime scales with estimating the gaze for multiple people per image (Fig. 7(b)). We measure the inference time for 1-10 people per image for both our default method, and our variant that uses a head position token $(t_{\text{pos}})$ and decodes gaze via cross attention and a dot product with the scene features (Tab. 7 configuration 1b 2b). Because the majority of our model’s computation can be attributed to the DINOv2 scene encoder ($>$95% of computation), which is run once regardless of the number of people, our model’s runtime does not increase much with addition of more people (15ms for 1 person vs. 19ms for 10 people). The token variant of our model with cross attention scales even better, as it decodes gaze for all people from the same final feature map. However, as shown in Fig. 7(a), this is accompanied by a slight performance decrease.
@@
10 Comparison to ViTGaze
We acknowledge concurrent work ViTGaze [55], which also proposes a single-branch transformer architecture for gaze target estimation based on DINOv2 pretrained weights. In contrast to Gaze-LLE, ViTGaze fully trains its ViT-S backbone (initialized from DINOv2 weights) end-to-end, and uses the attention weights between image patches as its feature representation. Gaze-LLE has the advantage of using the frozen DINOv2 features out-of-the-box, which is ideal for settings where general-purpose features are pre-computed and used for several downstream tasks. With its smaller backbone, ViTGaze is lightweight and may be better suited for on-device applications, where Gaze-LLE’s ViT-B or ViT-L backbone may be too large to run. We note that ViTGaze produces predictions for the L2 metric differently than prior gaze methods: while prior work determines the maximal gaze point from a standard-sized $64\times 64$ heatmap, ViTGaze uses additional postprocessing  [72] to bypass the limitations of the low resolution of the output heatmap. ViTGaze also uses a higher input resolution (512).
@@ 11 Performance with Estimated Head Bounding Boxes
Tu et al. [66] report that 2-stream methods suffer major performance drops when using head bounding boxes from a detector rather than the dataset ground truth. In contrast, we observe almost no performance degradation when pairing our method with a YOLOv5 head detector trained on CrowdHuman [54, 1] (Tab. 15). This result demonstrates that our single-stream design, which uses a coarse, downsampled head position map, is less dependent on an exact head crop, and works well with out of the box head detections. Given DINOv2’s strong performance on tasks such as semantic segmentation with linear probing, future work may explore integrating head detection directly into the pipeline by predicting heads from the same frozen DINOv2 features.
@@ 14 LoRA Backbones
To investigate if training the backbone improves performance, we explore using Low Rank Adaptation (LoRA) [27] on GazeFollow in Tab. 19. LoRA updates the backbone while introducing limited additional learnable parameters by learning weight update matrices as low rank decompositions. We update the query and value projections of the DINOv2 backbone using rank 16. We observe limited improvements, which we attribute to (1) the effectiveness of the frozen encoder’s feature representation for our task and (2) that our models with frozen encoders already achieve extremely close performance to the inter-rater performance of the human annotators, which serves as a soft upper bound on the L2 metrics.